# A Detailed Factor Analysis for the Political Compass Test: Navigating Ideologies of Large Language Models

### v2

@@ Abstract
Political Compass Test (PCT) or similar questionnaires have long been used to quantify LLM’s political leanings. Building on a recent line of work that examines the validity of PCT tests, we demonstrate that variation in standard generation parameters does not significantly impact a model’s PCT scores. However, external factors such as prompt variations and fine-tuning individually and in combination affect the same. Finally, we demonstrate that when models are fine-tuned on text datasets with higher political content than others, the PCT scores are not differentially affected. We also generalize these findings to a similar popular test. This calls for a thorough investigation into the validity of PCT and similar tests, as well as the mechanism by which political leanings are encoded in LLMs.
@@
Sadia Kamal†, Lalu Prasad Yadav Prakash†, S M Rafiuddin†, Mohammed Rakib†, Arunkumar Bagavathi†, Atriya Sen†, Sagnik Ray Choudhury‡ †Oklahoma State University, ‡University of North Texas {sadia.kamal,lprakas,srafiud,mohammed.rakib,atriya.sen}@okstate.edu, b.arun410@gmail.com, sagnik.raychoudhury@unt.edu
@@ 1 Introduction
Language models are now incorporated into many aspects of information access, decision support, and content generation, and consequently, the “political bias” or “leanings” of these models is under scrutiny. Defining what counts as “political bias” is challenging, as, unlike factual queries, politically charged questions often have no single objectively correct answer. In practice, this is operationalized in various ways, including measuring alignment with a particular wing on the left–right spectrum (in the US Aldahoul et al. (2025) or globally Rettenberger et al. (2025)), alignment with specific parties or candidates Aldahoul et al. (2025), and skew on individual social issues McGee (2023).
@@
A large number of recent studies Feng et al. (2023); Motoki et al. (2024); He et al. (2024) use the Political Compass Test111https://www.politicalcompass.org or PCT, a collection of 62 multiple-choice questions, where the respondent must agree on a Likert Scale (strongly disagree to strongly agree). These responses are then aggregated 222The aggregation function is not public. to generate two distinct scores, a social score and an economic score, each ranging from to . LLMs are generally prompted with each statement (possibly phrased as a question), and their level of agreement is recorded to infer the ideological coordinates (Figure 1).
@@
Theoretical validity issues Faulborn et al. (2025) aside, PCT has been shown to suffer from empirical instability when used with LLMs. For example, Röttger et al. (2024) shows that the models’ answers flip when they are forced into the PCT’s multiple-choice format and change again with minimal paraphrases or in open-ended settings, revealing large prompt-sensitivity and low test–retest reliability. This problem has also been discussed in the context of framing bias Bang et al. (2024), where the authors proposed an alternate mechanism to understand models’ political bias. However, despite these criticisms, PCT is still used in recent papers Liu et al. (2025); Ye et al. (2025); Rozado (2024), and few studies have rigorously evaluated the internal and external factors that can affect an LLM’s text generation, and consequently, affect its PCT score. We bridge this gap by investigating two research questions:
@@
Which common decoding parameters, if any, affect PCT results? Decoding parameters do have a substantial effect on generations, but how that translates to the final PCT results is underexplored. One can illustrate how the scores change (Figure 1), but are these differences significant and systematic? We answer using ANOVA tests on five common LLMs with varying sizes and three standard decoding parameters and find that the number of beams significantly affects the PCT results for some of the models, but overall, these parameters don’t affect the scores much. However, the prompt variation, as expected Röttger et al. (2024), has strong effects (§3).
@@
How does fine-tuning affect PCT? This research question has two motivations. On the operational side, the parameter changes induced by fine-tuning naturally alter a model’s generation, but how that affects the PCT scores is unknown. Fine-tuning should not have any effect when controlled for prompt variations, as it introduces little information that can alter a model’s political leanings. However, we do find evidence of significant effects (for illustrative purposes, see Figures 2 and 3 in the appendix A.2). On the cognitive side, this raises a question of whether this could be attributed to the text on which the models are fine-tuned. Specifically, we use two types of fine-tuning datasets – ones that contain political text, and ones that don’t. Arguably, human political leanings can change in response to new information, and we hypothesize that fine-tuning serves as a good proxy for this process in the models. We create a large collection of PCT tests by fine-tuning multiple LLMs on eight datasets, but can not find a significant effect of the dataset type (§4).
@@
PCT is one among many benchmarks for measuring political leanings in humans that have been studied in the context of LLMs Rozado (2024). We reproduce our findings on “8 Values Test” IDRlabs (2023), another such popular benchmark, highlighting the generalizability of our work. Similar to the PCT, the 8 Values Test also degenerates into LLMs/humans producing four scores across four axes (by answering a set of questions, A.4), that we use as dependent variables in our analyses.
@@
Beyond the general concern about the validity of PCT and similar tests, our conclusions are: a) The LLM PCT scores are possibly robust against variation in the generation parameters, and b) Researchers should verify their conclusions against both base and fine-tuned models. We hope this study inspires further investigation into how fine-tuning changes the political leanings of LLMs.
@@ 2 Experiment Setup
We use four open-source LLMs: Llama3-8B-Instruct Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Falcon3-7B-Instruct Almazrouei et al. (2023), and Gemma-3-4b-it Team et al. (2025). These models are widely used for chat and instruction-based applications and are well-known for their instruction-following capabilities. 333We use the smaller versions of these models as we fine-tune them later, but previous work has not found the scale to be a determining factor for PCT scores either Röttger et al. (2024). Also, we use 4-bit quantized versions of these models. For all experiments, we prompt (eg. “Choose one of the following options”) the models with the PCT/8 Values test statements (eg., “I’d always support my country, whether it was right or wrong”) and generate responses that we post-process and send to the PCT/8 Values server, and get back the scores.
@@ 3 RQ1: Decoding Params & Prompting
Our first experiment is to investigate the effect of standard decoding parameters on the PCT/8 Values tests. We use the ten prompts described in Röttger et al. (2024), and for each prompt, we generate responses from the models by varying the following decoding parameters: top_k, temperature, and num_beams. top_k constrains the decoding probability space to the most important k tokens. A higher temperature value increases the variability of generation. A higher number of beams improves the quality at the possible cost of diversity. We choose values for each parameter.
We assume that these factors (and the prompts) should not have interaction effects (eg., the number of beams should not depend on the prompts or vice versa); therefore, we run one-way ANOVA tests using the social scores and economic scores as dependent variables (8 Values have equivalent variables, A.4) and the decoding parameters as the independent ones. We use Levene’s test Levene (1960) to determine if the group variances are equal, and use Welch’s one-way ANOVA test Welch (1951) (which re-normalizes the degrees of freedom) when they are not.
The results are presented in Tables 3 (PCT, A.3) and 7 (8 Values, A.4). For 8 Values, none of the parameters has a significant effect for any model, and for PCT, only num_beams has a significant impact in Falcon (p-value ).
Confirming prior work’s conclusions Röttger et al. (2024), we find that prompting has a significantly low p-value (i.e., strong effect) in Economic scores as well as a high F-statistic (Table 4, A.3). However, we do not see such significance across the board for Social values. For 8 Values (Table 8, A.4), however, the prompts significantly affect all dependent variables across all models.
@@ 4 RQ2: Fine-Tuning
Having established that the decoding parameters don’t have a significant effect on the PCT tests, our next goal is to analyze the impact of fine-tuning. We investigate a diverse set of four natural language processing tasks (a) Classification, b) Conversation, c) Question-Answering, and d) Summarization) and eight distinct datasets for fine-tuning. For each of these tasks, we fine-tune the models with a control and a target dataset. A control dataset has textual content that is supposed to be neutral, i.e., non-politically oriented, so it should not impact the PCT scores. The target datasets, on the other hand, have text with strong political connotations, which could affect the trained models’ PCT score. The details of the datasets used in the experiments are provided in A.1.
@@
The details of the training process are described in the appendix A.6, and the evaluation results are discussed in A.7. In essence, we utilize PEFT methods that modify the parameters of attention matrices and generate nine model instances for each model class (Llama3/Mistral, etc.). One instance is the base model, and the other eight are its fine-tuned versions on the eight datasets.444We do not produce multiple model instances for the same base model and fine-tuning dataset by varying the initialization process, as our experiments suggest they are functionally equivalent. We train three instances of each model class on the SciSumm dataset using different seeds, but their test results do not vary significantly as illustrated in Table 18. We produce the PCT scores for these models by varying the prompts and other parameters as before, yielding a total of PCT test results across the base and the fine-tuned models. 555Ideally, we should generate results ( model types, models of each type, combinations of decoding params, and prompts), but some finetuned models are unable to generate responses for some PCT questions, and hence, are discarded. This number is for 8 Values.
@@
First, we aim to determine if the process of fine-tuning itself has an impact on PCT/8 Values scores. We find that to be true – the average PCT scores on the social and economic axes (and for equivalent variables in 8 Values) differ significantly across the base vs fine-tuned versions of the models as measured by independent t-tests Virtanen et al. (2020). See Table 5, A.3 for PCT and Table 9, A.4 for 8 Values.
@@
However, it is expected that the PCT/8 Values score of the fine-tuned model will depend on the prompt, and we are interested in observing the effect of fine-tuning while considering the effect of prompts. Therefore, we use two-way ANOVA tests with two independent variables: a) a categorical variable recording the prompt variation, and b) a binary variable indicating whether the model was fine-tuned or not. We test for the homogeneity of variances (Levene’s test) and normality of residuals (Shapiro–Wilk test Shapiro and Wilk (1965)), and when these conditions are violated, we use the Aligned Rank Transformed (ART) ANOVA Wobbrock et al. (2011) that first adjusts (or aligns) the data, then applies average ranks, allowing standard ANOVA methods to be used afterward.
@@
Table 1 and Table 10 (A.4) show the results for PCT and 8 Values, respectively. Individually, both prompting and fine-tuning have significant effects, as does their interaction. Importantly, fine-tuning should not have any effect on a model’s political leanings, but that happens to be the case. We conclude that studies examining the validity of PCT and similar tests should also consider the scores from fine-tuned versions of the base models.
@@
We hypothesize that changes in PCT/8 Values scores stem from the finetuning data: control-trained models should align with the base model, while target-trained ones should differ. To test this, for each task, we compute the group mean differences between the PCT/8 Values scores for base models and models trained on target or control datasets using the Games-Howell test Games and Howell (1976), which accounts for heteroscedasticity in our data. The results are presented in Tables 6 (PCT) and 12 (8 Values). For example, for the classification task, Gemma fine-tuned on the control dataset shows only a marginal difference from the base model on the PCT Social Score (), while fine-tuning on the target dataset yields a highly significant shift ().
@@
Tables 2 (PCT) and 11 (8 Values, A.4), derived from Tables 6 and 12, show the fraction of tasks where the finetuned model exhibits such a significant shift from the base model. Surprisingly, we observe that the distinction in content does not matter, i.e., the models change their scores independently of the content they are fine-tuned on. This creates an opportunity for exploring the mechanism by which finetuning changes the political leaning of LLMs, as measured by PCT and similar tests, which we leave for future work.
@@
Effect of model size & quantization. Given the computational cost of finetuning, we use one model size per family and its 4-bit quantized version. A natural question is whether the findings can be generalized to larger models and their non-quantized versions. To answer this, we repeat the PCT score experiments with Llama3.2-1B (a smaller variant of the Llama3-8B model used before) in both quantized and non-quantized forms. Tables 13, 14, 15, 16, 17 (A.5) present the results. Overall trends hold across sizes and quantization: decoding parameters have minimal impacts on PCT scores, fine-tuning leads to significant shifts (as measured by t-tests), and the effects of prompt and fine-tuning (and their interactions) are substantial. However, in contrast to Llama3-8B-quantized, the prompt variation does not significantly affect Economic scores in Llama3.2-1B-quantized. Otherwise, the results are consistent across different model sizes and quantized and non-quantized versions of the same size, supporting generalizability.

### v4

@@ Abstract
The Political Compass Test (PCT) and similar surveys are commonly used to assess political bias in auto-regressive LLMs. Our rigorous statistical experiments show that while changes to standard generation parameters have minimal effect on PCT scores, prompt phrasing and fine-tuning individually and together can significantly influence results. Interestingly, fine-tuning on politically rich vs. neutral datasets does not lead to different shifts in scores. We also generalize these findings to a similar popular test called 8 Values. Humans do not change their responses to questions when prompted differently (“answer this question” vs “state your opinion”), or after exposure to politically neutral text, such as mathematical formulae. But the fact that the models do so raises concerns about the validity of these tests for measuring model bias, and paves the way for deeper exploration into how political and social views are encoded in LLMs. The source code is publicly available here111https://github.com/sadiakamal/Detailed-factor-analysis-PCT.
@@
Sadia Kamal†, Lalu Prasad Yadav Prakash†, S M Rafiuddin†, Mohammed Rakib†, Atriya Sen†, Sagnik Ray Choudhury‡ †Oklahoma State University, ‡University of North Texas {sadia.kamal,lprakas,srafiud,mohammed.rakib,atriya.sen}@okstate.edu, sagnik.raychoudhury@unt.edu
@@ 1 Introduction
Language models are now incorporated into many aspects of information access, decision support, and content generation, and consequently, the political leanings of these models are under scrutiny. A large number of recent studies feng-etal-2023-pretraining; motoki2024more; he2024readingtweetsdecipheringideological measure models’ leanings through the Political Compass Test222https://www.politicalcompass.org or PCT, a collection of 62 multiple-choice questions, where the respondent must agree on a Likert Scale (strongly disagree to strongly agree). These responses are then aggregated 333The aggregation function is not public. to generate two distinct scores, a social score and an economic score, each ranging from to . LLMs are generally prompted with each statement (possibly phrased as a question), and their level of agreement is recorded to infer the ideological coordinates (Figure 1).
@@
Theoretical validity issues faulborn2025littlelefttheorygroundedmeasure aside, PCT has been shown to suffer from empirical instability when used with LLMs. For example, röttger2024politicalcompassspinningarrow shows that the models’ answers flip when they are forced into the PCT’s multiple-choice format and change again with minimal paraphrases for instructions to answer a question. This reveals a pattern of high prompt sensitivity and low test–retest reliability. However, despite these criticisms, PCT is still used in recent papers liu2025turning; ye2025large; llm-pol-preference, and few studies have rigorously evaluated the internal and external factors that can affect an LLM’s text generation, and consequently, affect its PCT score. We bridge this gap by investigating two research questions:
@@
Which common decoding parameters, if any, affect PCT results? Decoding parameters do have a substantial effect on generations, but how that translates to the final PCT results is underexplored. One can illustrate how the scores change (Figure 1), but are these differences significant and systematic?
We answer using ANOVA tests on four common LLMs with varying sizes and three standard decoding parameters and find that the number of beams significantly affects the PCT results for some of the models, but overall, these parameters have a minimal impact on the scores. However, the prompt variation has strong effects (§3) as expected röttger2024politicalcompassspinningarrow.
@@
How does fine-tuning affect PCT? This research question has two motivations. On the operational side, the parameter changes induced by fine-tuning naturally alter a model’s generation, but how that affects the PCT scores is unknown. Non-targeted fine-tuning should not affect PCT results when controlled for prompt variations, as it introduces little information that can alter a model’s leanings. However, we do find evidence of significant effects (for illustrative purposes, see Figures 2 and 3 in the appendix A.2). On the cognitive side, this raises a question of whether this could be attributed to the text on which the models are fine-tuned. Specifically, we use two types of fine-tuning datasets – those containing political text, and those that don’t. Arguably, human political leanings can change in response to new information, and we hypothesize that fine-tuning serves as a good proxy for this process in the models. We create a large collection of PCT tests by fine-tuning our LLMs on eight datasets, but can not find a significant effect of the dataset type (§4).
@@
PCT is one among many benchmarks for measuring political leanings in humans that have been studied in the context of LLMs llm-pol-preference. We reproduce our findings on “8 Values Test” 8values, another such popular benchmark, highlighting the generalizability of our work. Similar to the PCT, the 8 Values Test also degenerates into LLMs/humans producing four scores across four axes (by answering a set of questions, A.4), that we use as dependent variables in our analyses.
@@
This study raises concern about the validity of PCT and similar tests that anthropomorphize LLMs. We show that a) while the LLM PCT scores are possibly robust against variation in the generation parameters, they are significantly affected by fine-tuning and prompt variations, and b) the behavior of models as measured by these tests changes counter-intuitively when fine-tuned. We hope to inspire further investigation into the mechanism of how political leanings are encoded in LLM parameters.
@@ 2 Experiment Setup
We use four open-source LLMs: Llama3-8B-Instruct grattafiori2024llama, Mistral-7B-Instruct-v0.3 jiang2023mistral7b, Falcon3-7B-Instruct almazrouei2023falconseriesopenlanguage, and Gemma-3-4b-it gemmateam2025gemma3technicalreport. These models are widely used for chat and instruction-based applications and are well-known for their instruction-following capabilities. 444We use the smaller versions of these models as we fine-tune them later, but previous work has not found the scale to be a determining factor for PCT scores either röttger2024politicalcompassspinningarrow. Also, we use 4-bit quantized versions of these models. We discuss the effect of model size and quantization in A.5. For all experiments, we prompt (eg. “Choose one of the following options”) the models with the PCT/8 Values test statements (eg., “I’d always support my country, whether it was right or wrong”) and generate responses that we post-process and send to the PCT/8 Values server, and get back the scores.
@@ 3 RQ1: Decoding Params & Prompting
Our first experiment is to investigate the effect of standard decoding parameters on the PCT/8 Values tests. We use the ten prompts described in röttger2024politicalcompassspinningarrow, and for each prompt, we generate responses from the models by varying the following decoding parameters: top_k, temperature, and num_beams. top_k constrains the decoding probability space to the most important k tokens. A higher temperature value increases the variability of generation. A higher number of beams improves the quality at the possible cost of diversity. We choose values for each parameter. We aim to determine whether there is a statistically significant difference between the social and economic scores (for PCT, 8 Values have equivalent variables) in these results that can be attributed to these variations.
We assume that these factors (and the prompts) should not have interaction effects (eg., the number of beams should not depend on the prompts or vice versa); therefore, we run one-way ANOVA tests using the social scores and economic scores as dependent variables (8 Values have equivalent variables, A.4) and the decoding parameters as the independent ones. 555We use Levene’s test levene1960 to determine if the group variances are equal, and use Welch’s one-way ANOVA test welch1951 (which re-normalizes the degrees of freedom) when they are not.
The results are presented in Tables 5 (PCT, A.3) and 9 (8 Values, A.4). For 8 Values, none of the parameters has a significant effect for any model, and for PCT, only num_beams has a significant impact in Falcon (p-value ).
Confirming prior work’s conclusions röttger2024politicalcompassspinningarrow, we find that prompting has a significantly low p-value and a high F-statistic (Table 6, A.3) in Economic scores, i.e., has a strong effect. However, we do not see such significance across the board for Social values. For 8 Values (Table 10, A.4), however, the prompts significantly affect all dependent variables across all models.
@@ 4 RQ2: Fine-Tuning
Having established that the decoding parameters have no significant impact on the PCT test scores of LLMs, our next goal is to analyze the impact of fine-tuning. We investigate a diverse set of four natural language processing tasks: a) Classification, b) Conversation, c) Question-Answering, and d) Summarization, and eight datasets for fine-tuning. For each of these tasks, we fine-tune the models with a control and a target dataset. A control dataset has textual content that is supposed to be neutral, i.e., non-politically oriented, so it should not impact the PCT scores. The target datasets, on the other hand, have text with strong political connotations, which could affect the trained models’ PCT score. The details of the datasets used in the experiments are provided in A.1.
@@
The details of the training process are described in the appendix A.7, and the fine-tuning evaluation results are discussed in A.8. In essence, we utilize PEFT methods that modify the parameters of attention matrices and generate nine model instances for each model class (Llama3/Mistral, etc.). One instance is the base model, and the other eight are its fine-tuned versions on the eight datasets. In all datasets, the fine-tuned model performs better than the base one. 666We do not produce multiple model instances for the same base model and fine-tuning dataset by varying the initialization process, as our experiments suggest they are functionally equivalent. We train three instances of each model class on the SciSumm dataset using different seeds, but their test results do not vary significantly as illustrated in Table 25. We produce the PCT scores for these models by varying the prompts and other parameters as before, yielding a total of PCT test results across the base and the fine-tuned models. 777Ideally, we should generate results ( model types, models of each type, combinations of decoding params, and prompts), but if a finetuned model can’t generate responses for PCT questions, eg., generates unrelated text, we discard the entire test. This number is for 8 Values.
@@
First, we aim to determine if the process of fine-tuning itself affects PCT/8 Values scores. We find that to be true – the average PCT scores on the social and economic axes (and for equivalent variables in 8 Values) differ significantly across the base vs fine-tuned versions of the models as measured by independent t-tests 2020SciPy-ttest. See Table 7, A.3 for PCT and Table 11, A.4 for 8 Values.
@@
However, it is expected that the PCT/8 Values score of the fine-tuned model will depend on the prompt, and we are interested in observing the effect of fine-tuning while considering the effect of prompts. Therefore, we use two-way ANOVA tests with two independent variables: a) a categorical variable recording the prompt variation, and b) a binary variable indicating whether the model was fine-tuned or not. 888We test for the homogeneity of variances (Levene’s test) and normality of residuals (Shapiro–Wilk test shapiro1965), and when these conditions are violated, we use the Aligned Rank Transformed (ART) ANOVA art that first adjusts (or aligns) the data, then applies average ranks, allowing standard ANOVA methods to be used afterward.
@@
Table 1 and Table 12 (A.4) show the results for PCT and 8 Values, respectively. Individually, both prompting and fine-tuning have significant effects, as does their interaction. Importantly, fine-tuning, especially through PEFT, should not affect a model’s political leanings because it introduces minimal parameter changes, yet we find that to be the case.
@@
We hypothesize that changes in PCT/8 Values scores can stem from the finetuning data: control-trained models should have similar PCT scores as the base model, while target-trained ones could differ. To test this, for each task, we compute the group mean differences between the PCT/8 Values scores for base models and models trained on target or control datasets using the Games-Howell test games-howell, which accounts for heteroscedasticity in our data. The results are presented in Tables 8 (PCT, §A.3) and 14 (8 Values, §A.4). For example, for the classification task, Gemma fine-tuned on the control dataset shows only a marginal difference from the base model on the PCT Social Score (, assuming can be rejected at the significance level of ), while fine-tuning on the target dataset yields a highly significant shift ().
@@
Tables 2 (PCT) and 13 (8 Values, §A.4), derived from Tables 8 and 14, show the fraction of tasks where the finetuned model exhibits such a significant () shift from the base model. Surprisingly, we observe that the distinction in content does not matter, i.e., the models change their scores independently of the content they are fine-tuned on. This creates an opportunity for exploring the mechanism by which finetuning changes the political leaning of LLMs, as measured by PCT and similar tests, which we leave for future work.
@@
Model shift analysis: For each PCT/8 values question, the model should answer: 1. Strongly Agree, 2. Agree, 3. Disagree, or 4. Strongly Disagree. A qualified annotator evaluated each question to determine if an “agree” response aligned the model with the left or right of the social or political spectrum. For instance, agreement with “No one chooses their country of birth, so it’s foolish to be proud of it” indicates a left-leaning view, while agreement with “I’d always support my country, right or wrong” indicates a right-leaning one. After fine-tuning or prompt variation, a model can make four types of moves. If an original “Strongly Agree” (left) response changes to “Agree”, it’s a standard right move; if it changes to “Disagree” or “Strongly Disagree”, it’s a strong right move (see §A.6 for details).
Each model can make up to moves — computed as questions (prompt decoding variations) fine-tuning datasets. Fine-tuning has a substantial impact: with all else constant, models change their answers in a large proportion of cases (Llama: , Mistral: , Gemma: , Falcon: ), particularly Llama and Mistral, which is consistent with our statistical analysis.
Does the fine-tuning dataset influence the number of moves? We would expect control datasets to cause some movement, and target datasets to cause significantly more. However, as shown in Table 3 for Llama, this pattern does not hold universally. Both target and control fine-tunings produce a substantial number of moves. While target datasets cause more moves than controls in classification and summarization, this is not the case for conversation and QA tasks. The absence of this expected target–control distinction is consistent across all models and aligns with our statistical analysis.
| Task | Dataset name | %ge of move |
| Classification | IMDB (control) | 36.71 |
| Newsarticles (target) | 44.38 | |
| QA | OpenR1 (control) | 47.71 |
| CanadianQA (target) | 46.49 | |
| Conversation | Finetome (control) | 50.60 |
| Pol-convo (target) | 41.67 | |
| Summarization | Scisumm (control) | 28.77 |
| Newsroom (target) | 38.14 |
| Task | Dataset | Left | Right | ||||
| Standard | Strong | Total | Standard | Strong | Total | ||
| Classification | IMDB | 54.72 | 9.88 | 64.60 | 24.46 | 10.94 | 35.40 |
| Newsarticles | 38.86 | 7.07 | 45.93 | 28.88 | 25.18 | 54.07 | |
| QA | OpenR1 | 37.42 | 9.56 | 46.98 | 13.63 | 39.39 | 53.02 |
| CanadianQA | 57.16 | 15.47 | 72.63 | 9.01 | 18.37 | 27.37 | |
| Conversation | Finetome | 16.36 | 11.19 | 27.55 | 54.46 | 17.99 | 72.45 |
| Pol-convo | 64.78 | 9.43 | 74.21 | 11.30 | 14.49 | 25.79 | |
| Summarization | Scisumm | 50.43 | 11.10 | 61.53 | 21.76 | 16.71 | 38.47 |
| Newsroom | 38.70 | 16.15 | 54.85 | 22.06 | 23.09 | 45.15 | |
We next examine whether move strength (standard vs. strong) depends on the fine-tuning dataset. Table 4 reports the percentage of standard, strong, and total left/right moves for Llama on the PCT test. Two main patterns emerge: a) when moving rightward, Llama tends to make strong moves more often than standard ones, whereas leftward moves are predominantly standard; and b) control datasets generally induce movement opposite to that of the target datasets. These patterns, however, are model-dependent. As shown in Table 22 (Appendix), Falcon (which exhibits the fewest moves overall) shifts primarily leftward, with strong right movements being almost negligible. Additional results for 8 Values are presented in Tables 23 and 24 in §A.6.
This lack of clear patterns in the fine-grained analysis aligns with our aggregate statistical results and leads to two conclusions. First, fine-tuning experiments cast doubt on the validity of PCT and similar survey-based tests. Second, we need to better understand the mechanisms by which fine-tuning alters models’ encoding of political leaning.
Effect of model size & quantization. Given the computational cost of finetuning, we use one model size per family and its 4-bit quantized version. A natural question is whether the findings can be generalized to larger models and their non-quantized versions. To answer this, we repeat the PCT score experiments with Llama3.2-1B (a smaller variant of the Llama3-8B model used before) in both quantized and non-quantized forms. Tables 15, 16, 17, 18, 19 (A.5) present the results. Overall trends hold across sizes and quantization: decoding parameters have minimal impacts on PCT scores, fine-tuning leads to significant shifts (as measured by t-tests), and the effects of prompt and fine-tuning (and their interactions) are substantial. However, in contrast to Llama3-8B-quantized, the prompt variation does not significantly affect Economic scores in Llama3.2-1B-quantized. Otherwise, the results are consistent across different model sizes and quantized and non-quantized versions of the same size, supporting generalizability.